### v2

# Segment anything model 2: an application to 2D and 3D medical images

Abstract
Segment Anything Model (SAM) has gained significant attention because of its ability to segment various objects in images given a prompt. The recently developed SAM 2 has extended this ability to video inputs. This opens an opportunity to apply SAM to 3D images, one of the fundamental tasks in the medical imaging field. In this paper, We extensively evaluate SAM 2’s ability to segment both 2D and 3D medical images by first collecting 18 medical imaging datasets, including common 3D modalities such as computed tomography (CT), magnetic resonance imaging (MRI), and positron emission tomography (PET) as well as 2D modalities such as X-ray and ultrasound. Two evaluation pipelines of SAM 2 are considered: (1) multi-frame 3D segmentation, where prompts are provided to one or multiple slice(s) selected from the volume, and (2) single-frame 2D segmentation, where prompts are provided to each slice. The former is only applicable to 3D modalities, while the latter applies to both 2D and 3D modalities. Our results show that SAM 2 exhibits similar performance as SAM under single-frame 2D segmentation, and has variable performance under multi-frame 3D segmentation depending on the choices of slices to annotate, the direction of the propagation, the predictions utilized during the propagation, etc.
1 Introduction
Medical image segmentation is a crucial task for multiple clinical applications such as disease diagnosis and clinical analysis [17, 24, 27, 33]. Despite the advancements in medical imaging technologies, segmentation remains a challenging task due to the labor-intensive nature of data annotation and the complexity of medical images [4, 21, 36].
Segment Anything Model (SAM) addresses these challenges in multiple directions. On the one hand, SAM has demonstrated impressive zero-shot segmentation performance with prompt inputs, significantly reducing the need for extensive manual data annotation [6, 21]. On the other hand, several works fine-tune SAM to specific tasks and demonstrate improvements over standard segmentation techniques [10, 7, 18, 23, 19], such as nn-UNet [14]. Despite these advancements, SAM’s limitation to 2D images restricts its applicability to scenarios that require three-dimensional understanding [28].
Existing work addresses this challenge by introducing extra components to SAM to enable its 3D segmentation capability. For example, SAM3D [5] combines the SAM encoder with a lightweight 3D CNN decoder; 3DSAM-A [9] modifies the original prompt encoder and mask decoder to operate in 3D; SAM-Med3D [34] introduces an additional 3D convolution before the image encoder and replaces 2D positional encoding layers with 3D one. On the other hand, the recently introduced SAM 2 [28] solves this limitation fundamentally by extending the backbone of SAM to 3D. Specifically, SAM 2 proposes a memory bank that retains information from past predictions and allows it to make predictions on slices without prompts based on the information. This feature motivates us to examine SAM 2’s ability to segment 3D medical images since video segmentation can be transferred to 3D segmentation seamlessly, i.e., each slice can be treated as a frame. Note that we will use the terms “slice” and “frame” interchangeably throughout the paper.
In this paper, we extend the previous evaluation experiments on SAM [21] to SAM 2, aiming to explore the model’s effectiveness in a more complex, three-dimensional context. Specifically, we consider two evaluation pipelines: multi-frame 3D segmentation and single-frame 2D segmentation. In single-frame 2D segmentation [21], prompts are provided to the object of interest on each slice, whereas in multi-frame 3D segmentation, prompts are provided on one or a few slices selected from the volume. We further consider some unique challenges in multi-frame 3D segmentation, such as the selection of the slice(s) to be annotated, the direction of propagation, i.e., predicting slices without prompts based on ones with prompts, and the selection of prediction during propagation. Our experiments are conducted across 18 datasets, covering 5 modalities (magnetic resonance imaging (MRI), computed tomography (CT), positron emission tomography (PET), X-ray, and Ultrasound) and various body locations. We observe several trends of SAM 2 under both evaluation pipelines and summarize them in Section 5.
2 Methods
Figure 1: The pipeline of evaluating SAM 2 in the 3D setting. Different modes at each stage are proposed and evaluated.
In this section, we discuss the two evaluation pipelines, single-frame 2D segmentation and multi-frame 3D segmentation, in detail. Intersection over Union (IoU) is the evaluation metric throughout the paper. To have a comparable performance between 2D and 3D segmentation, IoU is only computed over non-empty slices.
2.1 Evaluation of Single-Frame 2D Model Testing on 2D Images
During single-frame 2D segmentation, SAM 2 exhibits the same behavior as SAM in segmenting the object of interest based on prompts. For datasets with 2D modalities, we run SAM 2 naturally at the image level. For datasets with 3D modalities, we simulate prompts for every slice of the volume. Following previous work [21], we design single-frame 2D segmentation in a non-iterative manner in which all prompts are determined without feedback from any prior predictions. Specifically, the following four 2D prompting modes (P-Mode) are used:
  1. 1.
    P-Mode 1: One point prompt placed at the center of the largest connected region of the object of interest.
  2. 2.
    P-Mode 2: One point prompt placed at each separate connected region of the object (up to three points).
  3. 3.
    P-Mode 3: One box prompt placed at the center of the largest connected region of the object of interest.
  4. 4.
    P-Mode 4: One box prompt placed at each separate connected region of the object (up to three boxes).
These strategies cover the most efficient and straightforward annotation strategy that provides one point or box prompt to each connected region of the object of interest.
Figure 2: Examples from all 18 datasets, each overlaid with annotation masks. The top three rows feature 12 examples from 3D datasets, while the bottom row presents 6 examples from 2D datasets.
2.2 Evaluation of Multi-Frame 3D Model Testing on 3D Volumes
SAM 2 differs from SAM mostly in its new ability to segment videos, which can be seamlessly transferred to the 3D image segmentation task. In this section, we are mostly interested in SAM 2’s semi-supervised segmentation ability, where we only provide prompts for one or a few frames in advance and use SAM 2 to predict other frames.
Initial Frame Selection. In video segmentation, the initial frame to be annotated is typically the first frame of video as it arrives first in the time stream and the object of interest usually does not change shape or size dramatically between consecutive frames [35][25]. This is also the strategy used by SAM 2 for the semi-supervised video object segmentation task. However, such a strategy may be ineffective in 3D medical imaging because, in most scenarios, the boundary of the object of interest appears first, and the change between different frames can be significant. Instead, when annotating medical volumes, annotators tend to start with the slice with the object of interest being the most salient. To evaluate SAM 2’s original frame selection (choosing the first slice) and to approximate human behavior in annotation medical volumes, we consider the following three frame modes (F-Modes):
  1. 1.
    F-Mode 1: Choose the edge slice when the object of interest first appears.
  2. 2.
    F-Mode 2: Choose the slice at the center of the object of interest.
  3. 3.
    F-Mode 3: Uniformly select 3 slices from the entire object of interest.
Prompts Simulation. When the slice to be annotated is selected, we only utilize the same point-based strategies (P-Mode 1 and 2) for simplicity because SAM 2’s video predictor does not have the function of adding boxes yet.
Abbreviated dataset name Full dataset name and citation Modality Num. classes Object(s) of interest Num. masks
Xray-Chest Montgomery County and Shenzhen Chest X-ray Datasets [15] X-ray 1 Chest 704
Xray-Hip X-ray Images of the Hip Joints [12] X-ray 2 Ilium, Femur 140
US-Breast Dataset of Breast Ultrasound Images [1] Ultrasound 1 Breast 630
US-Kidney CT2US for Kidney Segmentation [32] Ultrasound 1 Kidney 4,586
US-Muscle Transverse Musculoskeletal Ultrasound Image Segmentations [20] Ultrasound 1 Muscle 4,044
US-Ovarian-Tumor Multi-Modality Ovarian Tumor Ultrasound (MMOTU) [37] Ultrasound 1 Ovarian tumor 1,469
Table 1: 2D datasets evaluated in this paper: “num. masks” refers to the number of images with non-zero masks.
Propagation Direction. When the first frame is annotated, SAM 2 supports propagation to predict other non-conditioned frames (frames without prompts). This propagation step involves the newly introduced memory structures: memory attention, memory encoder, and memory bank. Memory attention conditions the current slice’s features on features from previously predicted slices and prompts. The memory encoder compresses these predictions into smaller embeddings and stores this information in the memory bank.
Note that since new prediction relies on the output from existing predictions, the order of propagation matters. Although the order of propagation might have less impact on video predictions in which the consecutive frames are similar and the object of interest appears consistently from beginning to end, it can have a more significant impact on 3D medical volumes in which the change between slices can be large and the object of interest does not occupy the full volume. To take the order of propagation into account, we consider the following 2 direction modes (D-Mode):
  1. 1.
    D-Mode1: Front-to-End Propagation: start from the first slice of the volume and propagate forward through the volume.
  2. 2.
    D-Mode2: Start from the first annotated slice and propagate backward to the first slice; then restart from the annotated slice and propagate forward to the end of the volume.
Predicted Mask Selection. Due to the ambiguity of providing click prompts, it is common to predict various levels of objects based on a single or a few clicks. For example, if we place a point prompt on a tumor inside the brain, the object of interest can either be the tumor or the whole brain.
In the 2D setting of both SAM and SAM 2, the mask predictor can create multiple mask outputs from which humans can select the one closest to their needs. If there is no human involvement in the selection step, we can simply pick the first prediction channel, which, according to previous SAM empirical results, tends to represent the smallest area among the three channels [21]. Alternatively, we can select the channel with the highest confidence based on the IoU calculated by SAM’s mask decoder.
In 3D mode, to prevent propagating ambiguity across frames, SAM 2 chooses to pick the prediction with the highest confidence and only utilizes this prediction when estimating the object of interest on other frames. However, based on previous experiments with SAM [21, 11], selecting the predictions with the highest confidence can be sub-par when compared to selecting the first channel in the medical imaging field. This observation motivates us with the following predicted mask selection strategies:
  1. 1.
    Selecting the predicted mask with the largest estimated IoU, i.e., confidence.
  2. 2.
    Selecting the first channel.
In summary, under multi-frame 3D segmentation, we cover 3 slice selection modes, 2 prompt-simulation modes, 2 predicted mask strategies, and 2 propagation modes, resulting in a total of 24 different experimental designs. Notably, we do not enumerate all possible choices. For some options, when we identify one choice that consistently performs better than the rest, we will trim combinations with other choices to reduce the total number of experiments.
Abbreviated dataset name Full dataset name and citation Modality Num. classes Object(s) of interest Num. masks Num. volumes
MRI-Spine Spinal Cord Grey Matter Segmentation Challenge [26] MRI 2 Gray matter and spinal cord 551 40
MRI-Heart Medical Segmentation Decathlon [31] MRI 1 Heart 1,301 20
MRI-Prostate Initiative for Collaborative Computer Vision Benchmarking [16] MRI 1 Prostate 1854 115
MRI-Brain The Multimodal Brain Tumor Image Segmentation Benchmark (BraTS) [22] MRI 3 GD-enhancing tumor, Peritumoral edema, necrotic and non-enhancing tumor core 12,591 206
MRI-Breast Duke Breast Cancer MRI: Breast + FGT Segmentation[30, 13] MRI 2 Vessel and fibroglandular tissue 14438 100
CT-Colon Medical Segmentation Decathlon [31] CT 1 Colon cancer primaries 1,285 126
CT-HepaticVessel Medical Segmentation Decathlon [31] CT 1 Vessels 13,046 303
CT-Pancreas Medical Segmentation Decathlon [31] CT 1 Parenchyma and mass 8,792 281
CT-Spleen Medical Segmentation Decathlon [31]) CT 1 Spleen 1,051 41
CT-Liver The Liver Tumor Segmentation Benchmark (LiTS) [3] CT 1 Liver 5,501 131
CT-Organ CT Volumes with Multiple Organ Segmentations (CT-ORG) [29] CT 5 Liver, bladder, lungs, kidney, and bone 4,776 10
PET-Whole-Body A FDG-PET/CT dataset with annotated tumor lesions [8] PET/CT 1 Lesion 1,015 42
Table 2: 3D datasets evaluated in this paper: “num. masks” refers to the number of images with non-zero masks. For 2D segmentation models, slices are used as inputs.
3 Dataset
Consistent with the previous experimental study on SAM [21], this study utilizes 18 diverse medical datasets to evaluate the performance of SAM 2. Specifically, all datasets are evaluated during single-frame 2D segmentation, and twelve datasets with 3D modalities (MRI, CT, and PET) are evaluated during multi-frame 3D segmentation. The “US-Nerve” dataset[2], which was included in the previous SAM evaluation paper[21], has been removed from this project due to licensing regulations. We also updated the MRI-Breast dataset to include more annotated volumes. The pre-processing steps are consistent with the previous study, except that for SAM 2 3D mode, we have to convert the input images to JPG format instead of PNG. We also provided visual representations of the annotations masks for each dataset in Fig. 2.
3.1 2D Datasets
We included 6 2D datasets, 2 X-rays, and 4 ultrasounds, covering 7 different anatomical objects. Specifically, the X-ray datasets cover chest and hip joint segmentation, and the ultrasound datasets encompass a broader range of regions, including breast, kidney, muscle, and ovarian tumor segmentation. Detailed information on these datasets can be found in Table 1. To keep the format consistent with 3D datasets, the images of 2D datasets are also converted to JPG using the same pipeline.
3.2 3D Datasets
We included 12 3D datasets, 5 MRI, 6 CT, and 1 PET-CT, covering 20 different anatomical objects. The structures of the datasets are modified to fit the input requirements of SAM 2 3D, where each individual volume has its own folder, and each slice is numbered according to its position within the volume. Table 2 demonstrates more detailed information.
Figure 3: The performance of SAM 2 under single-frame 2D segmentation on 2D datasets. Four prompt modes are considered, with results ranked in descending order based on P-Mode 4.
Figure 4: The performance of SAM 2 under single-frame 2D segmentation on 3D datasets. Four prompt modes are considered, with results ranked in descending order based on P-Mode 4.
4 Experimental Results
4.1 Results of SAM 2 under Singe-frame 2D Segmentation
In this section, we present the performance of SAM 2 under single-frame 2D segmentation scenario across four prompting modes (refer to Section 2.1). Fig. 3 and 4 show the results for 2D and 3D datasets, respectively. The findings indicate that SAM 2’s single-frame 2D segmentation capability is comparable with that of SAM (a comparison of single-frame 2D segmentation results between SAM and SAM 2 is provided in Fig. 5). Similar to SAM, the performance of SAM 2 varies significantly across different datasets. For instance, SAM 2 achieves an impressive IoU of 0.907 on the Xray-Hip dataset for ilium but performs poorly with an IoU of 0.277 on the MRI-Spine dataset for gray matter.
Comparing the performance for different prompting modes, we find box prompts consistently provide better results than point prompts. Moreover, providing more points does not always improve the performance. One reason can be that although having more points provides extra information on the object of interest, it encourages SAM 2 to segment the larger object. For example, having points at discontinuous regions on the breast makes SAM 2 more likely to segment the entire breast than to segment vessels precisely.
Figure 5: The single-frame 2D segmentation performance of SAM under 4 prompting modes 24 segmentation tasks (in gray) and the difference between the performance of SAM 2 and SAM. The differences are highlighted in red (when SAM has a higher IoU) and green (when SAM 2 has a higher IoU)
4.2 Results of SAM 2 under Multi-frame 3D Segmentation
In this section, we investigate the impact of each component during multi-frame 3D segmentation.
Figure 6: Comparison of different mask prediction selection strategies. The difference is highlighted in green (when the first channel strategy has a higher IoU) and red (when the largest confidence strategy has a higher IoU)
4.2.1 Impact of Predicted Mask Selection
We investigate the impact of prediction selection strategies across two initial frame selection modes: F-Mode 2 and F-Mode 3, all within the single point click setting (P-Mode 1) and using the front-to-end propagation direction (D-Mode 1). The results, depicted in Fig. 6, demonstrate that for the annotated slice, opting for the first channel’s prediction achieves performance comparable to or better than SAM 2’s default method of choosing the most confident slice. Notably, for the MRI-Spine: SC dataset, selecting the first channel leads to an approximate 0.2 and 0.23 improvement in IoU when starting from the center slice and 3 uniform slices respectively.
These results suggest that selecting the first channel from the annotated slices is a better choice when compared to the highest confidence selection. Consequently, we will focus on this selection method in the following sections.
Figure 7: Comparison of results for different frame selection choices, different click prompt choices on D-mode 1: Front to End Propagation.
Figure 8: Comparison of results for different frame selection choices, different click prompt choices on D-mode 2: Bidirectional Propagation. Additionally, we include a new mode P-GT which provides the ground truth mask to the first selected slices.
4.2.2 Impact of Initial Frame Selection
As shown in Fig. 7, we compare the results of providing prompts on the edge slice, the center slice, and three uniformly distributed slices for front-to-end propagation. The result illustrates a clear trend where the performance improves from having one annotated frame (F-Mode 1 and 2) to having 3 frames (F-Mode 3). Specifically, when giving a single prompt (P-Mode 1), selecting the first and center slice yields an average IoU of 0.23 and 0.22 across all datasets, respectively, and selecting three uniformly distributed slices further improves the performance to 0.32 IoU.
4.2.3 Impact of Prompt Modes
For different prompt modes (as illustrated by the light blue bars compared to the dark blue bars in Fig. 7 and Fig. 8), we find that providing more positive prompts (P-mode 2) on the initial slices results in poorer performance compared to providing just a single point prompt (P-mode 1). Providing one or more prompts yields consistent results for most datasets. One explanation is that additional prompts are only provided when there is more than one disconnected region. In most medical image datasets, the target object usually appears as a single connected region. For datasets with multiple objects, such as the CT-Organ: Lung task, supplying more positive prompts can assist the model in accurately segmenting both sides of the lung, thereby enhancing performance. Conversely, for smaller targets like those in the MRI-Brain dataset, providing additional prompts may cause the model to over-segment, potentially including the entire brain, which can lead to more significant error accumulation during propagation and be less effective than segmenting only a part of it with fewer prompts. Additionally, when comparing the results of three uniform slices (F-Mode 3) and three-point clicks in one slice (P-Mode 2, F-Mode 1 or 2), we find that distributing a total of three click prompts across multiple slices as initial input is more effective than placing them all on the same slice.
4.2.4 Impact of Propagation Mode
The comparison between Fig. 7 and Fig. 8sity demonstrate the superior of bi-directional propagation. Namely, the average IoU is improved by 0.14 and 0.06 IoU for F-Mode 2 and 3 respectively. Note that we do not include F-Mode 1 in this scenario because only propagating from the edge toward the objects makes sense.
These findings suggest that, although SAM 2 supports prompting on any frame as a condition to predict new frames, the propagation is more effective when first predicting adjacent slices. The effectiveness of bidirectional propagation also shows that reverse propagation, i.e., predicting the (N-1)th slice based on the Nth slice, works well.
To further investigate prediction changes during propagation, we propose an oracle mode where the ground truth mask is provided to the selected slice(s). This allows us to disentangle the error from imperfect prediction of the prompted slices and those during propagation; it also mimics one usage where users annotate the current frame and want SAM 2 to predict the rest. The results are shown as the green bars (P-GT) in Fig. 8. The average IoU for F-Mode 2 (one center slice mode) is 0.49, while the average IoU for F-Mode 3 (three uniform slices mode) is 0.56. Despite there being consistent improvements over using predicted masks, using ground truth masks does not yield high a IoU on average, compared with the IoU of 0.78 reported in SAM 2’s video dataset. We hypothesize that there are two reasons for this behavior: (1) SAM 2 was trained on a frame width of 8 frames, and thus the memory attention is most effective when the propagated frame is within 8 frames of the annotated frame. Medical volumes range from 10 to a few hundred slices. Consequentially, the performance will be significantly affected when the volumes are deeper. (2) SAM 2’s primary objective is to segment any videos. Although this function can be transferred seamlessly between video segmentation and 3D segmentation, videos and medical volumes can differ significantly. Namely, the change between consecutive slices can be large for medical volumes due to the nature of medical imaging technology. In this case, we believe that using SAM 2 for 3D medical imaging segmentation requires further improvements.
Figure 9: Summary of the averaged performance (2D+3D datasets for the first subplot and 3D for others) for different prompting choices. F-Mode 1: center slice, P-Mode 1: single click, and D-Mode 2: Bidirectional is the optimal setting when only one prompt is provided for the entire volume. When we compare a specific mode, the other two modes remain the same as this optimal setting.
4.2.5 Comparison between Single-frame and Multi-frame Evaluation
Based on our previous experiments, we find that the optimal configuration when conducting multi-frame 3D segmentation is F-Mode2, D-Mode 2 with using the first channel during propagation. Under this setting, providing 1 prompt to the entire volume (P-Mode 1) gives an average performance of 0.36, whereas providing 1 prompt to each slice (SAM 2 under single-frame 2D segmentation) gives an average performance of 0.50. A similar trend is observed when providing 3 prompts to the slice (P-Mode 2), with the performance being 0.38 and 0.52 for single-frame and multi-frame evaluation. It is important to note that despite the differences in results, we cannot conclude that single-frame 2D segmentation is better than multi-frame 3D segmentation. This is because, in the single-frame setting, prompts are provided on every slice, whereas in the multi-frame setting, only up to three prompts are provided to the entire volume and SAM 2 utilizes the memory bank for predicting slices without prompts. We hope this discussion can inspire future work, such as more effectively utilizing the memory bank for medical image features, to narrow the gap between the two settings and even exceed the performance of 2D single-frame segmentation.
5 Conclusion
In this work, we conduct an initial investigation of the performance of SAM 2 in the medical imaging field. In addition to evaluating the 2D segmentation task, we investigate its ability to perform the 3D segmentation task, thanks to SAM 2’s ability to segment videos. Our findings indicate that under single-frame 2D segmentation,
  1. 1.
    SAM 2 exhibits similar performance to that of SAM.
Fig. 9 summarizes our observation under multi-frame 3D segmentation:
  1. 1.
    For the multi-mask outputs, selecting the first channel is better than selecting the channel with the largest confidence during propagation.
  2. 2.
    For the initial frame selection, selecting multiple slices is better than selecting one slice with the cost of providing more prompts, and selecting the center slice tends to be the most cost-effective choice.
  3. 3.
    For prompt selection, generally using one point prompt placed at the center of the largest connected region of the object of interest) improves accuracy compared to using one point prompt placed at each separate connected region of the object.
  4. 4.
    Bidirectional propagation, starting from the annotated slice, is a more effective strategy when compared to propagating from beginning to end.
  5. 5.
    The best average performance of SAM 2 in the 3D segmentation is 0.36 IoU when only one point prompt is provided to the entire volume; the performance will be improved to 0.50 IoU when one point prompt is provided to every slice of the volume.